# Einstein Fields: A Neural Perspective to Computational General Relativity

### v1

@@ Abstract
We introduce Einstein Fields, a neural representation that is designed to compress computationally intensive four-dimensional numerical relativity simulations into compact implicit neural network weights. By modeling the metric, which is the core tensor field of general relativity, Einstein Fields enable the derivation of physical quantities via automatic differentiation. However, unlike conventional neural fields (e.g., signed distance, occupancy, or radiance fields), Einstein Fields are Neural Tensor Fields with the key difference that when encoding the spacetime geometry of general relativity into neural field representations, dynamics emerge naturally as a byproduct. Einstein Fields show remarkable potential, including continuum modeling of 4D spacetime, mesh-agnosticity, storage efficiency, derivative accuracy, and ease of use. We address these challenges across several canonical test beds of general relativity and release an open source JAX-based library, paving the way for more scalable and expressive approaches to numerical relativity. Code is made available at https://github.com/AndreiB137/EinFields.
@@ 1 Introduction
Figure 1: A conceptual overview of EinFields training and downstream pipeline. (i) Premise: The Einstein field equations (EFEs) in Equation (1) are highly non-linear partial differential equations defined on a 4D spacetime manifold, describing the geometric nature of gravitation. Their solutions define the metric tensor field $g_{\alpha\beta}(x^{\mu})$, which encodes the full spacetime geometry and serves as a tensorial generalization of the gravitational potential. In this work, we parametrize $g_{\alpha\beta}(x^{\mu})$ using a neural network. (ii) Training: The training is conducted on the metric tensor fields defined on a 4D spacetime grid. EinFields instead fit a continuous signal on these discrete representations, thus modeling 4D spacetime as a continuum, and returning the metric tensor field for a 4D spacetime query coordinate $p\equiv(t,x)\in\mathscr{M}$ at arbitrary resolution. (iii) Sobolev supervision: The reconstruction quality of the metric and its derivatives is improved by augmenting Sobolev losses, i.e., metric Jacobian (neighborhood structure) and Hessian (curvatures). (iv) Validation and downstream tasks: Sobolev improved EinFields’ AD-based derivatives enable accurate point-wise retrieval of differential geometric quantities, such as the Levi-Civita connection (covariant derivative), geodesics, curvature tensors, and their invariants.
General relativity (GR) is a foundational theory of modern physics that describes gravity as the curvature of spacetime caused by mass and energy. This curvature, along with angles, distances, geodesics, and causality, is encoded in the metric. At the heart of GR lie the Einstein field equations (EFEs), a set of coupled, tensor-valued, non-linear hyperbolic-elliptic partial differential equations (PDEs) that govern the metric behavior. EFEs are exceptionally complex, and exact analytical solutions exist only in idealized cases.
To solve EFEs in more general settings, numerical relativity (NR) has emerged as a powerful computational tool, whose predictive accuracy has repeatedly proven to be remarkable. Notable successes include the high-precision modeling of astrophysical phenomena such as mergers of binary black holes (Abbott et al., 2016a; b; c), binary neutron stars (Hayashi et al., 2025), and neutron star–black hole systems (Abbott et al., 2017). These numerical models have also played a central role in the detection of gravitational waves (GWs) by the Laser Interferometer Gravitational-Wave Observatory (LIGO) (Collaboration et al., 2015) and Virgo (Acernese et al., 2014), leading to groundbreaking, Nobel-prize winning discoveries  (Castelvecchi, 2017). Likewise, the precision of applied technologies, most notably the global positioning system (GPS), also depends on subtle relativistic effects, such as gravitational time dilation, which constitutes the influence of gravity on the flow of time. Incorporating these corrections and recalibrating the GPS satellites is crucial for providing accurate real-time positioning data, as even tiny relativistic discrepancies lead to significant errors in location estimates.
Nonetheless, state-of-the-art NR is one of the most computationally intensive domains of scientific computing, typically requiring petascale computing infrastructures (Huerta et al., 2019). With a new generation of more sensitive detectors underway, e.g., the Einstein Telescope (Punturo et al., 2010; Abac et al., 2025) and the Laser Interferometer Space Antenna (Amaro-Seoane et al., 2017), there is a strong demand for further improving the performance and accuracy of NR.
With the growing success of machine learning-based approaches to scientific modeling, particularly neural PDE surrogates for spatiotemporal dynamics and forecasting (Thuerey et al., 2021; Zhang et al., 2023; Brunton et al., 2020; Li et al., 2021; Brandstetter et al., 2022; Bodnar et al., 2025; Brandstetter, 2025), and neural fields (NeFs) for compact representations of scenes (Mildenhall et al., 2021), shapes (Park et al., 2019; Chen & Zhang, 2019; Mescheder et al., 2019), images (Karras et al., 2021), audio, video (Sitzmann et al., 2020), or physical quantities (Raissi et al., 2019), it is worth exploring the potential for hybrid methodologies that integrate machine-learning and traditional NR frameworks. In the context of GR, we hypothesize that NeFs offer promising advantages for modeling continuum fields, where data are costly to store and accurate differentiation is crucial for capturing physical laws. However, the key difference to conventional NeFs is that when encoding the spacetime geometry of GR into NeFs representations, dynamics emerge naturally as a byproduct.
In this work, we introduce Einstein Fields or, for short, EinFields – a representation that can parameterize computationally intensive four-dimensional NR simulations as compact neural network weights. A schematic of EinFields is shown in Figure 1. Contrary to traditional NR approaches, EinFields model the spacetime and the tensor-valued fields defined on it as a continuum. EinFields are NeFs that parametrize the metric tensor field $(t,x)\in\mathbb{R}^{1,3}\mapsto g_{\alpha\beta}(t,x)\in\mathbb{R}^{4\times 4}$ (or $\mathbb{R}^{10}$ due to symmetry). Since the metric and its first two derivatives entirely encode the geometric structure of the spacetime and, thus, many physical laws and phenomena, we leverage automatic-differentiation (AD) that are native to NeFs  (Griewank, 2003; Griewank & Walther, 2008; Baydin et al., 2018). Since many physical phenomena predicted by GR, such as the precession of orbits, the bending of light, or GW formation, require high precision when modeled computationally, achieving small errors is a central quality criterion for EinFields. We show the impact of the model, the optimizer, the loss function, impact of the coordinate system, as well as the differentiation strategy, successfully achieving an agreement down to $\mathtt{1E{-8}}$ relative error and recovering these sensitive physical phenomena in simulated experiments.
It is worth stressing that this work is designed to encode existing simulations into neural network weights, rather than serving as a surrogate solver.
Unlike NeFs in existing applications, e.g., for representing 3D scenes or geometries, EinFields must respect the non-trivial aspects of differential geometry and gravitational physics: (i) encapsulate all features of a (pseudo-) Riemannian metric, such as raising and lowering of contravariant and covariant indices of other tensors; (ii) capture the underlying physics, not coordinates-associated artifacts; (iii) yield highly accurate metric tensor derivatives, pivotal for reconstructing dynamics, such as geodesics (Equation (2)), curvature invariants (Equation (79)), and satisfying the EFEs (Equation (1)).
Owing to the properties of NeFs, EinFields inherit the following advantages:
@@
  • Neural compression. EinFields encode the geometric information in compact neural network representations (typically less than 2 million parameters), agreeing with ground-truth metric tensor components up to $1\mathtt{E{-9}}$ relative precision. These models offer highly memory-efficient approximations of complex geometric structures.
  • Discretization-free. The continuity of NeFs enables the modeling of gravitational events without the discretization of the domain. These neural implicits can be fit with arbitrary point samples (e.g., regular grids, irregular grids, unstructured point clouds) and can be evaluated at arbitrary points of the spacetime.
    This avoids resolution-limited artifacts and increases flexibility and ease of use.
  • Tensor differentiation. EinFields are smooth ($C^{\infty}(\mathcal{M})$) MLP parameterizations. This enables querying higher-order tensor derivatives, e.g., Christoffel symbols, Riemann tensor, and curvature invariants, via point-wise AD.
    We reconstruct differential geometric quantities with AD, which is accurate and easy to use in contrast to higher-order finite-difference (FD) methodologies that require adaptive mesh refinement (AMR)  (Berger & Oliger, 1984) near high-curvature regions.
Our secondary set of contributions mainly concerns the training and validation-related aspects:
  • Reconstructing seminal tests of GR. We use synthetic data generated from analytical solutions to validate and characterize EinFields. Beyond conventional ML evaluation metrics, we use the reparameterized metrics to evaluate derived physical quantities and reconstruct famous experimental predictions of GR. These include (i) perihelion precession around Schwarzschild metrics; (ii) geodesics around Kerr metrics with varying spins and ring-singularity through Kretschmann invariant; (iii) deformation of a ring due to gravitational waves, Weyl scalar $\Psi_{4}$, and radiated power of outgoing fields in the linearized gravity setting. These phenomena are due to small corrections from Newtonian gravity and serve as sensitive indicators of the metric quality.
  • JAX-based differential geometry library. We introduce a differential geometry library based on JAX (Bradbury et al., 2018), leveraging its AD and accelerated linear algebra (XLA) features (just-in-time compilation and automatic vectorization). Our library includes:
    1. 1.
      Tensor algebra: handling of co- and contravariant indices, tensor contractions, and specialized operations;
    2. 2.
      Tensor calculus: a suite of differential geometric quantities, operations (Lie derivatives, covariant derivatives), and identities common in GR;
    3. 3.
      Collection of common analytical solutions of EFEs.
@@ 2 Background
Our work combines two primary ideas: general relativity from physics and neural fields from machine learning. We begin with a brief introduction of GR, stressing three key properties that pertain to our work: GR is a field theory, GR is coordinate independent, and physics is entirely contained in the metric and its first two derivatives. For this, we shall introduce the technical concepts of tensors and their coordinate transformations. For a more complete introduction, we refer to Appendix A and the sources therein.
2.1 General Relativity
2.1.0.1 GR is a tensor field theory.
GR, formulated by Albert Einstein, extends Newtonian gravity with a geometric interpretation of gravity, where mass and energy tell spacetime how to curve, and curved spacetime tells objects how to move (Misner et al., 2017). This is formalized by Einstein’s field equations (EFEs)
@@ 2.1 General Relativity
EFEs are a set of 10 coupled non-linear, tensor-valued, second-order partial differential equations and can be viewed as a tensorial generalization of the Newton-Poisson equation for gravity $\nabla^{2}\Phi(r)=-4\pi G\rho(r)$ (Misner et al., 2017; Poisson, 2004). In EFEs, $G_{\alpha\beta}=R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}$ is the Einstein tensor, formed from the metric tensor $g_{\alpha\beta}$, the Ricci curvature tensor $R_{\alpha\beta}$, and the Ricci curvature scalar $R$. Thus, the left-hand side of EFEs is entirely described by the metric and its derivatives, with $\Lambda$ being the cosmological constant. The right-hand side depends on the stress-energy tensor $T_{\alpha\beta}$ describing the matter distribution, with $G$ being Newton’s constant.
The backbone of these highly complicated field equations are tensors, and the mathematical formalism is differential geometry. Spacetime is modeled as a four-dimensional continuum with the structure of a pseudo-Riemannian (Lorentzian) manifold $\mathcal{M}$. A pseudo-Riemannian manifold is equipped with an indefinite metric that allows for both positive and negative distances.
2.1.0.2 Tensors.
A rank ($r$, $s$) tensor $T$ at a point $x\in\mathcal{M}$ is the multilinear map from $r$ covectors and $s$ vectors to a real number:
$T:\underbrace{\mathcal{V}^{*}\times...\times\mathcal{V}^{*}}_{r-\text{copies}}\times\underbrace{\mathcal{V}\times...\times\mathcal{V}}_{s-\text{copies}}\rightarrow\mathbb{R}\ .$
The $r$ vectors and $s$ covectors pair with the respective $r$ covariant and $s$ contravariant components of the tensor. As such, a tensor is an element that lives in a tensor product of vector and dual spaces, i.e., $T\in(\mathcal{V})^{\otimes r}\otimes(\mathcal{V}^{*})^{\otimes s}$.
A tensor in a particular choice of basis $\{e_{\alpha_{n}}\}_{1\leq n\leq r}\in\mathcal{V}$ and $\{\vartheta^{\beta_{n}}\}_{1\leq n\leq s}\in\mathcal{V}^{*}$ is given by
$T=T^{\alpha_{1}\alpha_{2}\dots\alpha_{r}}_{\;\;\;\;\;\;\;\;\beta_{1}\beta_{2}\dots\beta_{s}}e_{\alpha_{1}}\otimes\cdots\otimes e_{\alpha_{r}}\otimes\vartheta^{\beta_{1}}\otimes\cdots\otimes\vartheta^{\beta_{s}}\ ,$
where $T^{\alpha_{1}\alpha_{2}\dots\alpha_{r}}_{\;\;\;\;\;\;\;\;\beta_{1}\beta_{2}\dots\beta_{s}}\equiv T(\vartheta^{\alpha_{1}},\cdots,\vartheta^{\alpha_{r}},e_{\beta_{s}},\cdots,e_{\beta_{s}})$ are the components of the tensor in this particular basis.
Importantly, the abstract tensor itself is a coordinate-independent object. Similarly, EFEs are diffeomorphism covariant, meaning physical laws remain unchanged under differentiable coordinate transformations (Misner et al., 2017; Tao, 2008).
2.1.0.3 Tensor fields.
A tensor field is a collection of tensor-valued quantities such that at each point $x\in\mathcal{M}$, the multilinear function associates a value $T_{x}\in\mathcal{V}_{x}^{\otimes p}\otimes\big{(}\mathcal{V}^{*}\big{)}_{x}^{\otimes q}$. Consequently, the components $T^{\alpha_{1}\ldots\alpha_{r}}_{\beta_{1}\ldots\beta_{s}}(x)$ are functions of the points of the manifold. For example, both the metric tensor field $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$, i.e., $g=g_{\alpha\beta}(x)\ dx^{\alpha}\otimes dx^{\beta}$ and Ricci tensor $R_{\alpha\beta}(x)$ are of rank $(0,2)$. The Ricci curvature scalar $R$ is a rank 00 tensor: $R:\mathcal{M}\rightarrow\mathbb{R}$, and, similarly, velocity and gradient vector fields constitute rank $(1,0)$ and rank $(0,1)$ tensors, respectively [^1] Colloquially speaking, the components of a vector transform contravariantly, i.e., opposite to how the base set transform, and the components of a contravariant vector transform covariantly, i.e., in the same way as the base set transforms..
2.1.0.4 Metric.
The metric is a rank $(0,2)$ symmetric bilinear form $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ that generalizes the notion of an inner product on the tangent space $T_{x}\mathcal{M}$ of a differentiable manifold $\mathcal{M}$ (Jost, 2008). It enables the computation of angles between vector fields defined on the manifold and provides a means to measure distances via the line element, expressed in local coordinates as $ds^{2}=g_{\alpha\beta}(x^{\mu})\,dx^{\alpha}dx^{\beta}$, where $g_{\alpha\beta}$ denotes the components of the metric tensor in a chosen coordinate system, which may be curvilinear and need not be intrinsically Euclidean (see Appendix A.3.2 for detailed exposition). In relativistic physics, the metric is often denoted as $ds^{2}$ or $d\tau^{2}$, and in general relativity, it serves as the tensorial generalization of the “gravitational potential”, encoding the causal structure and geometric properties of spacetime. The spacetime metric tensor $g_{\alpha\beta}$ is a $4\times 4$ symmetric matrix with ten independent components, and together with its first and second derivatives, fully determines a wide range of physical phenomena. These include gravitational wave propagation, gravitational lensing, time dilation, black hole dynamics, and the evolution of astrophysical systems under the influence of gravity.
2.1.0.5 Derivative operators.
2.1.0.6 Differential geometric objects.

### v2

@@ Abstract
We introduce Einstein Fields, a neural representation designed to compress computationally intensive four-dimensional numerical relativity simulations into compact implicit neural network weights. By modeling the metric, the core tensor field of general relativity, Einstein Fields enable the derivation of physical quantities via automatic differentiation. Unlike conventional neural fields (e.g., signed distance, occupancy, or radiance fields), Einstein Fields fall into the class of Neural Tensor Fields with the key difference that, when encoding the spacetime geometry into neural field representations, dynamics emerge naturally as a byproduct. Our novel implicit approach demonstrates remarkable potential, including continuum modeling of four-dimensional spacetime, mesh-agnosticity, storage efficiency, derivative accuracy, and ease of use. It achieves up to a $\mathtt{4,000}$-fold reduction in storage memory compared to discrete representations while retaining a numerical accuracy of five to seven decimal places. Moreover, in single precision, differentiation of the Einstein Fields-parameterized metric tensor is up to five orders of magnitude more accurate compared to naive finite differencing methods. We demonstrate these properties on several canonical test beds of general relativity and numerical relativity simulation data, while also releasing an open-source JAX-based library: https://github.com/AndreiB137/EinFields, taking the first steps to studying the potential of machine learning in numerical relativity.
@@ 1 Introduction
General relativity (GR) describes gravity as the curvature of four-dimensional spacetime, encoded in the metric tensor and governed by the Einstein field equations (EFEs), a system of coupled, nonlinear hyperbolic-elliptic PDEs. Exact solutions are available only for idealized cases, so numerical relativity (NR) has become essential for accurate modeling of astrophysical events. Notable successes of NR include the high-precision modeling of black hole mergers (Abbott et al., 2016a; b; c), high-precision binary neutron star merger simulations (Hayashi et al., 2025), and neutron star–black hole systems (Abbott et al., 2017). NR has also been central to the confirmation of gravitational waves (GWs) detected by LIGO and Virgo interferometers, leading to Nobel-prize–winning discoveries.
Nonetheless, state-of-the-art NR is one of the most computationally intensive domains of scientific computing, requiring massive parallelization on petascale computing infrastructures (Lovelace, 2021; Huerta et al., 2019). This is due to several computational challenges in NR, including adaptive high-resolution spatial discretization and finite-difference (FD) methods on these, which, in addition, are vulnerable to numerical errors in sensitive regions. Moreover, NR simulations are equally storage-intensive, producing up to petabytes of data per run, prohibiting the storage and distribution of simulations on HPC systems (Reed & Dongarra, 2015).
Recent progress in machine learning for scientific computing has shown the potential of hybrid neural-classic workflows (Thuerey et al., 2021; Zhang et al., 2023; Brunton et al., 2020; Li et al., 2021; Brandstetter et al., 2022; Bodnar et al., 2025; Brandstetter, 2025). In addition, neural fields (NeFs) (Park et al., 2019; Müller et al., 2022; Chen & Zhang, 2019) have emerged as a powerful tool in visual computing for compact and continuous representations of traditionally discrete data, such as images, shapes, and physical fields, with ease of querying and differentiating. This raises the question of whether such hybrid approaches can advance next-generation computational GR workflows, particularly in handling and compressing tensorial quantities and their derivatives defined on adaptive high-fidelity discretizations.
To this end, we propose and investigate EinFields, which provide the following contributions:
@@
  • Neural compression. EinFields encode geometric information in compact neural representations with typically fewer than two million parameters. They reproduce metric tensor components with relative precision up to seven decimal digits (and up to nine in favourable coordinate charts). This yields memory-efficient approximations of complex spacetime geometries, with compression factors up to $\mathtt{4000\times}$ across analytical and numerical solutions.
  • Discretization-free representation. EinFields are trained on arbitrary point samples, including regular, irregular, and unstructured sets. They provide continuous query access to tensor fields at any resolution by learning these as continuous functions from discrete samples, which removes discretization artefacts.
  • Enhanced tensor differentiation. As smooth neural functions, EinFields support continuous evaluation of higher-order geometric quantities such as Christoffel symbols, Riemann tensors, and curvature invariants via point-wise automatic differentiation (Griewank & Walther, 2008). Initial results suggest that this approach can outperform high-order finite-difference methods on uniform grids, with accuracy gains up to $\mathbf{10^{5}}$ in FLOAT32.
  • High-fidelity reconstruction of tests of General Relativity. We evaluate the physical fidelity of EinFields on analytical GR solutions and assess derived observables in addition to standard ML metrics. The models faithfully reproduce key relativistic phenomena, including orbital precession in Schwarzschild and Kerr spacetimes, and allow accurate extraction of gravitational-wave distortions and strain. We further test a BSSN (Baumgarte & Shapiro, 2021) evolution of an oscillating neutron star. Early results indicate that EinFields scale to realistic numerical relativity workflows despite the complexity of the solution.
@@ 2 Background
Our work lies at the intersection of two domains: GR, along with its computational framework of NR, and NeFs, a ubiquitous tool from computer vision. While a complete introduction to GR and its mathematical backbone, differential geometry, is beyond the scope of this Section (see detailed exposition in Appendix A or succinct version in Appendix D), we stress three key properties that pertain to our work: (i) GR is a field theory of tensor-valued quantities, (ii) GR is intrinsically coordinate-independent, and (iii) gravitational physics is entirely encapsulated in the metric and its first two derivatives.
Tensors. A rank ($r$, $s$) tensor $T$ at a point $x\in\mathcal{M}$ is the multilinear map from $r$ covectors and $s$ vectors to a real number:
$T:\underbrace{\mathcal{V}^{*}\times...\times\mathcal{V}^{*}}_{r-\text{copies}}\times\underbrace{\mathcal{V}\times...\times\mathcal{V}}_{s-\text{copies}}\rightarrow\mathbb{R}\ .$ (1)
The $r$ vectors and $s$ covectors pair with the respective $r$ covariant and $s$ contravariant components of the tensor. As such, a tensor is an element that lives in a tensor product of vector and dual spaces, i.e., $T\in(\mathcal{V})^{\otimes r}\otimes(\mathcal{V}^{*})^{\otimes s}$. A tensor in a particular choice of basis $\{e_{\alpha_{n}}\}_{1\leq n\leq r}\in\mathcal{V}$ and $\{\vartheta^{\beta_{n}}\}_{1\leq n\leq s}\in\mathcal{V}^{*}$ is given by $T=T^{\alpha_{1}\alpha_{2}\dots\alpha_{r}}_{\;\;\;\;\;\;\;\;\beta_{1}\beta_{2}\dots\beta_{s}}e_{\alpha_{1}}\otimes\cdots\otimes e_{\alpha_{r}}\otimes\vartheta^{\beta_{1}}\otimes\cdots\otimes\vartheta^{\beta_{s}}\ ,$ where $T^{\alpha_{1}\alpha_{2}\dots\alpha_{r}}_{\;\;\;\;\;\;\;\;\beta_{1}\beta_{2}\dots\beta_{s}}\equiv T(\vartheta^{\alpha_{1}},\cdots,\vartheta^{\alpha_{r}},e_{\beta_{s}},\cdots,e_{\beta_{s}})$ are the components of the tensor in this particular basis and transforms as shown in Eq. (20). A tensor field assigns to each point $x\in\mathcal{M}$ a multilinear map, i.e. a tensor, $T_{x}\in\mathcal{V}_{x}^{\otimes p}\otimes(\mathcal{V}^{*})_{x}^{\otimes q}$. In appropriate coordinates, its components $T^{\alpha_{1}\ldots\alpha_{r}}_{\beta_{1}\ldots\beta_{s}}(x)$ vary smoothly across the manifold.
General relativity extends Newtonian gravity with a geometric interpretation of gravity, where mass and energy tell spacetime how to curve, and curved spacetime tells objects how to move (Misner et al., 2017). This is formalized by the Einstein’s field equations (EFEs)
@@
EFEs are a set of 10 coupled non-linear, tensor-valued, second-order partial PDEs and can be viewed as a tensorial generalization of the Newton-Poisson equation for gravity $\nabla^{2}\Phi(r)=-4\pi G\rho(r)$ (Misner et al., 2017; Poisson, 2004). In EFEs, $G_{\alpha\beta}=R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}$ is the Einstein tensor, formed from the metric tensor field $g_{\alpha\beta}(x^{\mu})$, which are solutions of the EFEs and tensorial generalization of the gravitational potential $\Phi(\mathbf{x})$. The Ricci curvature tensor $R_{\alpha\beta}$, and the Ricci curvature scalar $R$ are related by second derivatives $g_{\alpha\beta}$. Thus, the left-hand side of EFEs is entirely described by the metric and its derivatives, with $\Lambda$ being the cosmological constant. The right-hand side depends on the stress-energy tensor $T_{\alpha\beta}$ describing the matter distribution, with $G$ being Newton’s constant.
Metric tensor and its derivatives. The metric tensor is a rank $(0,2)$ symmetric bilinear form $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ that generalizes the notion of an inner product on the tangent space $T_{x}\mathcal{M}$ of a differentiable manifold $\mathcal{M}$ (Jost, 2008). It enables the computation of angles between vector fields and a means to compute distances via the line element: $ds^{2}=g_{\alpha\beta}(x^{\mu})\,dx^{\alpha}dx^{\beta}$. In GR, the components $g_{\alpha\beta}$ in a particular coordinate system can be seen as a $4\times 4$ symmetric matrix with ten independent components. The metric defines the causal structure and contains all geometric information of spacetime. Importantly, its partial derivatives $\partial$ yield the Christoffel symbols $\Gamma_{\alpha\beta\gamma}(x^{\mu})$, which describe the notion of parallel transport and defines a covariant derivative operation $\nabla_{\alpha}=\partial_{\alpha}+\Gamma_{\alpha}$ (all detailed in Appendix A.3.3). In turn, the connection’s derivatives (i.e., metric second-derivatives) yield the Riemann curvature tensor $R^{\delta}_{\alpha\beta\gamma}(x^{\mu})$, which encodes tidal forces of gravity. The trace part of $R^{\delta}_{\alpha\beta\gamma}$ (index contraction w.r.t. metric: $\text{Tr}_{g}$ – see Eq. (44)) is the Ricci tensor $R_{\alpha\beta}$, also a symmetric rank $(0,2)$ tensor. Its subsequent contraction yields the Ricci scalar $R$ (all detailed in Appendix A.3.5). This can be summarized schematically as follows (a more detailed pictorial version is available in Figure 10:
$g_{\alpha\beta}\xrightarrow{\partial}\Gamma^{\gamma}_{\alpha\beta}\xrightarrow{\nabla}R^{\delta}_{\ \alpha\beta\gamma}\xrightarrow{\text{Tr}_{g}}R_{\alpha\beta}\xrightarrow{\text{Tr}_{g}}R.$ (3)
Figure 1: A conceptual overview of EinFields training and downstream pipeline. (i) Premise: The Einstein field equations (EFEs) in Eq. (2) are highly non-linear PDEs defined on a 4D spacetime manifold, describing the geometric nature of gravitation. Their solutions define the metric tensor field $g_{\alpha\beta}(x^{\mu})$, which encodes the full spacetime geometry and serves as a tensorial generalization of the gravitational potential. In this work, we parametrize $g_{\alpha\beta}(x^{\mu})$ using a neural network. (ii) Training: The training is conducted on the metric tensor fields defined on 4D spacetime points, such as uniform or hierarchical grids. EinFields instead fit a continuous signal on these discrete representations, thus modeling 4D spacetime as a continuum, and returning the metric tensor field for a 4D spacetime query coordinate $p\equiv(t,x)\in\mathscr{M}$ at arbitrary resolution. (iii) Sobolev supervision: The reconstruction quality of the metric and its derivatives is improved by augmenting Sobolev losses, i.e., metric Jacobian (neighborhood structure) and Hessian (curvatures). (iv) Validation and downstream tasks: Sobolev improved EinFields’ AD-based derivatives enable accurate point-wise retrieval of differential geometric quantities, such as the Levi-Civita connection (covariant derivative), geodesics, curvature tensors, and their invariants.
Higer-order methods.  FD methods with adaptive mesh refinement (AMR) (Berger & Oliger, 1984) have long underpinned tensor calculus in NR, discretizing space and time with high-order stencils (Appendix C). An $n$-th order stencil yields truncation errors of $\mathcal{O}(h^{n})$, where $h$ is the grid spacing. Widely used fourth- or sixth-order schemes improve accuracy but incur larger communication costs due to broader stencil footprints in parallelized settings. In contrast, modern NR increasingly opts for (pseudo-)spectral methods (Scheel et al., 2025), which represent fields globally through polynomial bases (Fornberg, 1996), yielding an efficiency of up to $\mathtt{1000{-}5000}\times$ faster on CPUs than FD approaches on GPUs at comparable accuracy (Rashti et al., 2025).
Neural fields (NeFs), also known as implicit neural representations (INR) or coordinate-based neural networks, are multi-layer perceptrons (MLP) using very specific activation functions that are memory-efficient, implicit, continuous, infinitely differentiable maps, capable of capturing high-fidelity detail across complex domains (Xie et al., 2021; Essakine et al., 2024). Some well known NeFs are SIREN with sinusoidal activations (Sitzmann et al., 2020), WIRE with Gabor wavelet activations (Saragadam et al., 2023) for example. These properties have motivated their primary development and adoption in computer vision domains for representation, generation, and inversion tasks. Considering scientific computing domains, NeFs, when integrated with physics-informed losses (e.g., constraints and conservation laws), can be used for solving forward and inverse problems, including spatiotemporal dynamics governed by PDEs. In these settings, NeFs effectively act as PDE solvers, more commonly referred to as physics-informed neural networks (PINNs) (Raissi et al., 2019; Karniadakis et al., 2021; Wang et al., 2025b).
3 Method – Parametrizing Tensor fields via EinFields
Consider the four-dimensional spacetime $(\mathscr{M},g)$ (a manifold equipped with a metric) corresponding to an exact or numerical solution to the EFEs [^1] From now on, we omit the cosmological constant term $\Lambda g_{\alpha\beta}$.: $G_{\alpha\beta}=8\pi G\ T_{\alpha\beta}$. An EinField models the 10 independent components of the metric tensor field as a compact NeF, ultimately mapping the spacetime coordinates $x\equiv(x^{0},x^{1},x^{2},x^{3})$ to the symmetric rank $(0,2)$ metric tensor field:
$\hat{g}:x\in\mathscr{M}\rightarrow g_{\alpha\beta}(x)\in\text{Sym}^{2}(T^{*}_{x}\mathscr{M})\ .$ (4)
We deploy an MLP $\hat{g}_{\theta}$ with parameters $\theta$, denoted $\hat{g}$ for simplicity, to over-fit on the ground truth tensor field. Methodologically, this enables directly compressing the entire geometric information into storage-cheap NN weights, yielding continuous access (different from the training points) at arbitrary resolution of the metric and its non-trivial tensor differentiation (e.g., for Lie derivative or covariant derivatives) information devoid of mesh (re)construction on curvilinear manifolds. This generalizes to an arbitrary rank $(r,s)$ tensor field $T^{\alpha_{1}\ldots\alpha_{r}}_{\beta_{1}\ldots\beta_{s}}(x^{\mu})$. Thus, EinFields posit a neural alternative to address one or more of the challenges associated with traditional methods (typically utilizing higher-order finite differencing schemes) in NR by not relying on costly spatiotemporal discretizations. Our framework should be considered as a special case of Neural Tensor Fields.
3.1 Distortion
We define the distortion as the algebraic deviation of the spacetime metric from flat space,
$\Delta_{\alpha\beta}=g_{\alpha\beta}-\eta_{\alpha\beta}\ ,$ (5)
with the Minkowski background $\eta_{\alpha\beta}$ in that particular coordinate system. From a learning perspective, this decomposition acts as a preprocessing step that removes the offset (fixed background metric). Flat contributions, that may even dominate numerically (e.g., $g_{tt}\!\sim\!1/r$, $g_{\theta\theta}\!\sim\!r^{2}$), are removed, leaving only the non-trivial geometric content, such as the curvature. Thus, the network focuses its representational capacity on meaningful deviations rather than redundantly relearning flat-space structure. As we show in 4.2, this improves scaling, accelerates convergence, and emphasizes the dynamic, physically relevant features of the metric during training.
3.2 Retrieving physics via neural tensor field derivatives
Higher derivative losses. Beyond the metric tensor itself, its first- and second-order derivatives are critical to GR, as they govern geodesic motion, tidal forces, and curvature. Accurate trajectory reconstruction requires point-wise precise evaluation of Christoffel symbols and Riemann tensors. Such a high-fidelity extraction from EinFields is facilitated by Sobolev training (Czarnecki et al., 2017; Son et al., 2021), a formulation that explicitly incorporates higher derivative losses (Chetan et al., 2024; Wang et al., 2025b) – see Section F.4. The supervision on the metric Jacobian $\partial_{\mu}g_{\alpha\beta}$ (40 independent components) and Hessian $\partial_{\mu}\partial_{\nu}g_{\alpha\beta}$ (100 independent components) rectifies irregularities in the metric field and its derivatives, yielding substantial accuracy gains and consequentially improving the precision of point-wise Christoffel symbols $\Gamma^{\gamma}_{\alpha\beta}(x^{\mu})$ and Riemann tensors $R^{\sigma}_{\alpha\beta\gamma}(x^{\mu})$ queries by up to two orders of magnitude. This is given by the modified loss function:
$\mathcal{L}^{g}_{\text{Sob}}(\theta)=\mathbb{E}_{x}\bigg[\lambda_{0}\|g_{\alpha\beta}(x)-\hat{g}_{\alpha\beta}(x)\|^{2}+\sum_{j=1}^{2}\lambda_{j}\left\|\partial_{x}^{(j)}g_{\alpha\beta}(x)-\partial_{x}^{(j)}\hat{g}_{\alpha\beta}(x)\right\|^{2}\bigg]\ ,$ (6)
with $\lambda_{j}$ being some coefficients and $\partial_{x}^{(1)}\equiv\partial_{\mu}$ and $\partial_{x}^{(2)}\equiv\partial_{\nu}\partial_{\mu}$ written in a succinct notation. Instead of implementing higher-order FD stencils, our framework enables access to exact higher-order tensor derivatives via AD. This is illustrated in the AD workflow for differential geometry in Figure 2.